OpenStudy (anonymous):

4+9+14+19....+49. the sum of the sequence is?

OpenStudy (anonymous):

apply artmtic progrssn formula...

OpenStudy (anonymous):

Refer to the Mathematica 9 Home Edition attachment.

OpenStudy (akashdeepdeb):

This is an arithmetic progression with a common difference of 5 and having the first term = 4. a = 4 l = 49 d = 5 n = ? Find n using this formula: l = a + (n-1)d After finding 'n'. Use the formula for the sum of an arithmetic progression: S = (n/2) * [2a + (n-1)d]

OpenStudy (anonymous):

use \[l=a+(n-1)d\] to get n value number of terms will be 10 and use \[\sum=(n/2)*(a+l)\]for sum of sequence. answer will be 265

OpenStudy (k_lynn):

95